# Preserving Source Video Realism: High-Fidelity Face Swapping for Cinematic Quality

### v1

@@ 5.2 Comparisons with Existing Methods
In this section, we compare our model against several state-of-the-art face-swapping approaches, including SimSwap [chen2020simswap], InfoSwap [gao2021information], BlendSwap [shiohara2023blendface], CanonSwap [luo2025canonswap], DiffSwap [zhao2023diffswap], FaceAdapter [han2024face], our baseline model VACE [jiang2025vace], and the widely used industrial system Inswapper [facefusion2025], which is also employed for our keyframe generation and dataset construction. As shown in Tab. 2, on FF++—which mainly contains relatively simple scenarios such as interviews and livestreams—LivingSwap achieves the best overall ranking.
We further select the top-performing five methods on FF++ for additional evaluation on CineFaceBench. As shown in Tab. 1, LivingSwap achieves state-of-the-art performance across multiple metrics and average ranks on both the easy and hard identity settings of our CineFaceBench, demonstrating strong robustness under varying identity similarity levels. Compared to our keyframe generation system, Inswapper, although our keyframes are generated from its outputs, our model exhibits significantly superior temporal consistency, better preservation of source-video attributes, and more stable face-swapping behavior in difficult scenarios such as side views and occlusions (see Fig. 4). This also indicates that our approach is robust to imperfect or problematic keyframes, as further discussed in the Appendix C.
@@ 5.3 Ablation Studies
Ablation of Synthetic Data Quality. As discussed in Sec. 4.4, we construct a paired dataset, Face2Face, using a state-of-the-art single-frame face swapping model [facefusion2025] to investigate the impact of per-frame errors on data quality. In challenging scenarios, such models often introduce flickering artifacts, identity inconsistencies, and various visual distortions (see Fig. 5). On the other hand, these challenging cases also enhance the diversity of the dataset. To this end, we conduct a detailed analysis of how the distribution of simple versus complex scenarios affects overall data quality. Specifically, we filter the data pairs based on identity variation between the original and synthetic videos and categorize them into three groups for comparison: the entire dataset, the first 70% of the data, and the last 70% of the data. Experimental results presented in Tab. 3 show that neither ID-similar data nor data with significant identity discrepancies substantially improves the model’s identity performance. Instead, we observe that using the full dataset leads to better results in terms of lighting, pose. This suggests that the diversity of samples within the full dataset plays an indispensable role in enhancing model performance. Meanwhile, the results also demonstrate the model’s robustness to noisy training data, enabling it to ultimately go beyond the limitations of training data quality (Appendix A). Therefore, in the subsequent experiments, we adopt all available data pairs for model training.
@@
Ablation of Model Design. We conduct ablation studies on three key components of our model design: video reference, keyframe guidance, and target image reference. As shown in Tab. 4, when we replace the video reference with the traditional inpainting approach, the model exhibits a decline in fidelity metrics. Regarding identity injection, when we remove the keyframe guidance and rely solely on the target image, we observe a significant drop in identity similarity as well as a noticeable degradation in temporal consistency. Conversely, when we ablate the identity information provided by the target image, we still observe a decline in identity similarity. This is due to the limitations of keyframes in certain scenarios—such as occlusion, extreme angles, or closed eyes—which may result in the loss of critical identity features.
@@ 6 Conclusion
This work presented LivingSwap, the first video reference-guided face swapping model that leverages keyframes as conditioning signals to enhance both fidelity and temporal coherence in video face swapping. By combining keyframe conditioning with video reference guidance, our approach ensures stable identity preservation and high-fidelity reconstruction across long video sequences. We propose a novel paired dataset, Face2Face, along with a role-reversing strategy that provides reliable ground-truth supervision and tackles the challenge of scarce data for reference-guided training. Extensive experiments demonstrate that LivingSwap seamlessly integrating target identities with source video attributes such as expressions, lighting, and exhibiting strong performance in complex face-swapping scenarios. Our model significantly reduces manual effort in production workflows, enabling more efficient and flexible video editing in film and entertainment.
[…]
@@
  • Appendix D: Robustness to Identity Differences.
  • Appendix E: Robustness to Attribute Variations in Source Video.
  • Appendix F: Grayscale Keyframe Guidance for Robust Color Learning.
  • Appendix G: Face2Face Construction Details.
  • Appendix H: CineFaceBench Construction Details.
  • Appendix I: Comparison with Close-Source Methods.
Figure 6: Qualitative comparison between the data pairs in Face2Face (by Inswapper [facefusion2025]) and corresponding results generated by LivingSwap. Benefiting from reversing the role in data pair and strong priors in pretrained model, LivingSwap surpasses the quality of its training data, achieving better expression consistency and overall realism. Unlike Inswapper-based results, our method avoids local failure cases—such as incomplete swaps, mismatched regions, and occlusion-induced artifacts—demonstrating its strong generalization beyond the training dataset.
@@ Appendix A Generalization Beyond Train Data Quality
In Sec. 5.3, we analyzed the robustness of LivingSwap under varying levels of data quality, demonstrating the of our model to failed noisy train data. To further investigate whether LivingSwap is fundamentally constrained by the quality of the training dataset itself, we conduct an additional experiment. We manually select several noisy source–target pairs from Face2Face. These pairs contain various types of degradation, including local failure cases caused by failed swaps (e.g., residual beards), artifacts and misaligned expressions arising from large identity gaps or occlusions, as shown in Fig. 5. We then run LivingSwap on these noisy pairs using the exact same inference process as Inswapper, which was used to construct the dataset.
@@ Appendix B Keyframe Identity Injection for Accumulated Identity Errors
Figure 7: Keyframes Identity Injection for resolving Accumulated ID Errors. Keyframe Identity Injection for Resolving Accumulated ID Errors. When using only the first frame for ID injection, face-swapping results suffer from gradually accumulating ID errors over time. In contrast, with Keyframe Identity Injection, each video chunk is corrected individually by swapped keyframe, ensuring better ID consistency throughout the entire long video sequence.
@@ Appendix C Robustness to Keyframe Quality
Appendix D Robustness to Identity Differences
@@
For the scenario of swapping different identities for the same source video, we conducted experiments with multiple videos and identities. As shown in Fig. 9, leveraging the advantages of keyframe identity injection, LivingSwap achieves satisfactory results for the same video, regardless of whether the identity difference is large or small. We hypothesize that this robust of identity difference is due to the diversity of identities in our training data, as discussed in Sec. 5.3.
Appendix E Robustness to Attribute Variations in Source Video
@@
Appendix F Grayscale Keyframe Guidance for Robust Color Learning
@@
We finetune final checkpoint of LivingSwap in 5000 steps to adapt it to grayscale pipeline. As shown in Fig. 12, we observe that grayscale keyframe guidance significantly reduces color bleeding and flickering artifacts in challenging cases where the keyframe edits exhibit inconsistent or unrealistic colors, while maintaining comparable identity preservation and temporal consistency.
Appendix G Face2Face Construction Details
We construct our dataset Face2FaceSwap based on CelebV-Text [yu2023celebv] and VFHQ [xie2022vfhq]. First, we perform crop, resize, and clipping operations on the dataset to ensure the resolution is 640×640 pixels and the video length is approximately 200 frames. We then randomly pair the data and extract the first frame from the target video as the target face image. Next, we apply Inswapper [facefusion2025] to perform face-swapping on the entire dataset. The process is conducted using 8 NVIDIA H100 GPUs over a duration of 120 hours. Additionally, we use the face-parsing model [yu2018bisenet] to generate the face mask video. For the ablation study on the inpainting paradigm, we also use the pose estimation model [yang2023effective] to generate the corresponding pose video. After filtering out the failed samples from the processing steps, our dataset Face2Face contains a total of 152,221 video samples, with a cumulative duration exceeding 300 hours. Finally, we reverse the roles in each training pair: the swapped video is used as the source video (model input), while the original video serves as the target video (ground-truth supervision).
Appendix H CineFaceBench Construction Details
@@
Due to the fact that most scenes in FaceForensics++ [rossler2019faceforensics++] consist of relatively simple settings such as interviews, hosts, or live broadcasts, it does not evaluate critical aspects often required in film scenarios, such as the model’s ability to preserve facial expressions, lighting, makeup, and overall fidelity after face-swapping, as well as the stability of long video clips longer than 30 seconds. To address the limitations of the aforementioned benchmark, we have constructed a film scene face-swapping benchmark, CineFaceBench.
CineFaceBench consists of 200 video clips, paired with easy and hard target face images, resulting in 400 data pairs. Specifically, we downloaded and selected 100 video clips from two free video websites, Pixabay [pixabay] and Pexels [pexels]. Additionally, we selected 100 video clips from the OpenHumanVid dataset [li2025openhumanvid] and preprocessed them, resulting in 200 video samples used for evaluation. As shown in Fig. 13, these 200 clips include challenging examples from film scenes, featuring difficult scenarios such as unique lighting, exaggerated expressions, micro-expressions, special makeup, occlusions, and even facial deformations. In addition, there are several video cases that are longer than 1 minute.
@@
Appendix I Comparison with Close-Source Methods
Figure 14: Qualitative comparison with recent inpainting-based video face swapping methods [chen2024hifivfs, han2024face] shows that our approach better preserves source video attributes (e.g., lighting and expression) and achieves greater stability under occlusions.
Recently, several inpainting-based video face swapping methods using the Stable Video Diffusion model [blattmann2023stable] are proposed, such as HiFiVFS [chen2024hifivfs] and FaceAdapter [han2024face]. However, these methods are not open-source. To enable a comparison with them, we captured several demos from their project websites and conducted tests using the same target face image. The comparative results are shown in the Fig. 14. Our approach better preserves the original video attributes such as lighting and expression, and also demonstrates strong stability in occluded cases.

### v2

@@ 5.2 Comparisons with Existing Methods
In this section, we compare our model against several state-of-the-art face-swapping approaches, including SimSwap [4], InfoSwap [11], BlendSwap [33], CanonSwap [25], DiffSwap [52], FaceAdapter [14], our baseline model VACE [20], and the widely used industrial system Inswapper [15], which is also employed for our keyframe generation and dataset construction. As shown in Tab. 2, on FF++—which mainly contains relatively simple scenarios such as interviews and livestreams—LivingSwap achieves the best overall ranking.
We further select the top-performing five methods on FF++ for additional evaluation on CineFaceBench. As shown in Tab. 1, LivingSwap achieves state-of-the-art performance across multiple metrics and average ranks on both the easy and hard identity settings of our CineFaceBench, demonstrating strong robustness under varying identity similarity levels. Compared to our keyframe generation system, Inswapper, although our keyframes are generated from its outputs, our model exhibits superior temporal consistency, better preservation of source-video attributes, and more stable face-swapping behavior in difficult scenarios such as side views and occlusions (see Fig. 4). This also indicates that our approach is robust to imperfect or problematic keyframes, as further discussed in Supplementary Material.
@@ 5.3 Ablation Studies
Ablation of Synthetic Data Quality. We investigate the impact of synthetic data quality using Face2Face, constructed via Inswapper [15], which introduces flickering artifacts, identity inconsistencies, and various visual distortions (see Fig. 5). We first analyze the trade-off between quality and diversity by categorizing data into three groups based on identity variation (top 70%, bottom 70%, and full). Experimental results (Tab. 3) reveal that our model is robust to data quality variations, as filtering yields no benefit. Instead, the diversity inherent in the full dataset proves critical for enhancing overall fidelity, leading us to adopt all available data. To further verify whether LivingSwap is constrained by the quality of this training data, we evaluate it on manually selected noisy pairs containing artifacts. As illustrated in Fig. 6 in Supplementary Material, LivingSwap consistently surpasses the quality of the original training samples, producing results with improved expression alignment, visual realism, and significantly fewer local artifacts. We attribute this ability to generalize beyond noisy supervision to two key designs: (1) our data reversing strategy, which ensures reliable ground-truth supervision even when input pairs are noisy, and (2) strong pretrained priors, which robustly correct corrupted patterns.
@@
Ablation of Model Design. We conduct ablation studies on three key components of our model design: video reference, keyframe guidance, and target image reference. As shown in Tab. 4, when we replace the video reference with the traditional inpainting approach, the model exhibits a decline in fidelity metrics. Regarding identity injection, when we remove the keyframe guidance and rely solely on the target image, we observe a significant drop in identity similarity as well as a noticeable degradation in temporal consistency. Conversely, when we ablate the identity provided by the target image, we still observe a decline in identity similarity. This is due to the limitations of keyframes in certain scenarios, such as occlusion, extreme angles, or closed eyes, which may result in the loss of critical identity features.
@@ 6 Conclusion
This work presented LivingSwap, the first video reference guided face swapping model that leverages keyframes as conditioning signals to enhance both fidelity and temporal coherence in video face swapping. By combining keyframe conditioning with video reference guidance, our approach ensures stable identity preservation and high-fidelity reconstruction across long video sequences. We propose a novel paired dataset, Face2Face, along with a role-reversing strategy that provides reliable ground-truth supervision and tackles the challenge of scarce data for reference-guided training. Meanwhile, we construct a new benchmark, CineFaceBench, for the video face swapping task in cinematic scenes. Extensive experiments demonstrate that LivingSwap seamlessly integrates target identities with source video attributes while exhibiting strong performance. Our model significantly reduces manual effort in production workflows, enabling more efficient and flexible video editing in film and entertainment.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62576315, No. 62506338).
@@ Acknowledgments
  • Appendix D: Potential of Keyframe Selection.
  • Appendix E: Robustness to Identity Differences.
  • Appendix F: Robustness to Attribute Variations in Source Video.
  • Appendix G: Grayscale Keyframe Guidance for Robust Color Learning.
  • Appendix H: Face2Face Construction Details.
  • Appendix I: CineFaceBench Construction Details.
  • Appendix J: Comparison with Closed-Source Methods.
  • Appendix K: Limitations of LivingSwap.
Figure 6: Qualitative comparison between the data pairs in Face2Face (by Inswapper [15]) and corresponding results generated by LivingSwap. Benefiting from reversing the role in data pair and strong priors in pretrained model, LivingSwap surpasses the quality of its training data, achieving better expression consistency and overall realism. Unlike Inswapper-based results, our method avoids local failure cases—such as incomplete swaps, mismatched regions, and occlusion-induced artifacts—demonstrating its strong generalization beyond the training dataset.
@@ Appendix A Generalization Beyond Train Data Quality
In Sec. 5.3, we analyzed the robustness of LivingSwap under varying levels of data quality, demonstrating the robustness of our model to failed noisy train data. To further investigate whether LivingSwap is fundamentally constrained by the quality of the training dataset itself, we conduct an additional experiment. We manually select several noisy source–target pairs from Face2Face. These pairs contain various types of degradation, including local failure cases caused by failed swaps (e.g., residual beards), artifacts and misaligned expressions arising from large identity gaps or occlusions, as shown in Fig. 5. We then run LivingSwap on these noisy pairs using the exact same inference process as Inswapper, which was used to construct the dataset.
@@ Appendix B Keyframe Identity Injection for Accumulated Identity Errors
Figure 7: Keyframes Identity Injection for resolving Accumulated ID Errors. When using only the first frame for ID injection, face-swapping results suffer from gradually accumulating ID errors over time. In contrast, with Keyframe Identity Injection, each video chunk is corrected individually by swapped keyframe, ensuring better ID consistency throughout the entire long video sequence.
@@ Appendix C Robustness to Keyframe Quality
Appendix D Potential of Keyframe Selection
Table 5: The potential of rule-based keyframe selection.
Methods ID Sim. $\uparrow$ Expr. $\downarrow$ Light $\downarrow$ Gaze $\uparrow$ Pose $\downarrow$
Inswapper 0.224 2.609 0.248 0.587 3.867
LivingSwap (Fixed-interval Keyframe) 0.073 2.441 0.243 0.634 3.240
LivingSwap (Rule-based Keyframe) 0.240 2.393 0.236 0.681 3.504
Although LivingSwap demonstrates good robustness when handling keyframes with low face-swapping quality, using high-quality keyframes can significantly improve the overall video quality. Due to the limitations of current face-swapping models in handling profiles or extreme angles, we introduce a simple yet effective keyframe selection method: selecting frontal frames as keyframes based on face orientation detection (yaw within ±30° and pitch within ±20°). For evaluation, we selected the 10 worst-performing cases from the CineFaceBench for comparison. As shown in Tab. 5, using this straightforward keyframe selection rule greatly enhances the quality of the face-swapped video results.
Furthermore, by leveraging the keyframe guidance feature of LivingSwap, we can manually refine the face-swapping results or perform post-processing modifications (e.g., adjusting appearance or makeup) using tools like PhotoShop, providing greater flexibility in enhancing and fine-tuning the final face-swapping outcomes.
Appendix E Robustness to Identity Differences
@@
For the scenario of swapping different identities for the same source video, we conducted experiments with multiple videos and identities. As shown in Fig. 9, leveraging the advantages of keyframe identity injection, LivingSwap achieves satisfactory results for the same video, regardless of whether the identity difference is large or small. We hypothesize that this robustness of identity difference is due to the diversity of identities in our training data, as discussed in Sec. 5.3.
Appendix F Robustness to Attribute Variations in Source Video
@@
Appendix G Grayscale Keyframe Guidance for Robust Color Learning
@@
We finetune the final LivingSwap checkpoint for 5,000 steps to adapt it to grayscale pipeline. As shown in Fig. 12, we observe that grayscale keyframe guidance significantly reduces color bleeding and flickering artifacts in challenging cases where the keyframe edits exhibit inconsistent or unrealistic colors, while maintaining comparable identity preservation and temporal consistency.
Appendix H Face2Face Construction Details
We construct our dataset Face2Face based on CelebV-Text [47] and VFHQ [41]. First, we perform crop, resize, and clipping operations on the dataset to ensure the resolution is 640×640 pixels and the video length is approximately 200 frames. We then randomly pair the data and extract the first frame from the target video as the target face image. Next, we apply Inswapper [15] to perform face-swapping on the entire dataset. The process is conducted using 8 NVIDIA H100 GPUs over a duration of 120 hours. Additionally, we use the face-parsing model [46] to generate the face mask video. For the ablation study on the inpainting paradigm, we also use the pose estimation model [45] to generate the corresponding pose video. After filtering out the failed samples from the processing steps, our dataset Face2Face contains a total of 152,221 video samples, with a cumulative duration exceeding 300 hours. Finally, we reverse the roles in each training pair: the swapped video is used as the source video (model input), while the original video serves as the target video (ground-truth supervision).
Appendix I CineFaceBench Construction Details
@@
Due to the fact that most scenes in FaceForensics++ [30] consist of relatively simple settings such as interviews, hosts, or live broadcasts, it does not evaluate critical aspects often required in film scenarios, such as the model’s ability to preserve facial expressions, lighting, makeup, and overall fidelity after face-swapping, as well as the stability of long video clips longer than 30 seconds. To address the limitations of the aforementioned benchmark, we have constructed a film scene face-swapping benchmark, CineFaceBench.
CineFaceBench consists of 200 video clips, paired with easy and hard target face images, resulting in 400 data pairs. Specifically, we downloaded and selected 100 video clips from two free video websites, Pixabay [27] and Pexels [26]. Additionally, we selected 100 video clips from the OpenHumanVid dataset [23] and preprocessed them, resulting in 200 video samples used for evaluation. As shown in Fig. 13, these 200 clips include challenging examples from film scenes, featuring difficult scenarios such as unique lighting, exaggerated expressions, micro-expressions, special makeup, occlusions, and even facial deformations. In addition, there are several video cases that are longer than 1 minute.
@@
Appendix J Comparison with Closed-Source Methods
Figure 14: Qualitative comparison with recent inpainting-based video face swapping methods [5, 14] shows that our approach better preserves source video attributes (e.g., lighting and expression) and achieves greater stability under occlusions.
Recently, several inpainting-based video face swapping methods using the Stable Video Diffusion model [2] are proposed, such as HiFiVFS [5] and FaceAdapter [14]. However, these methods are not open-source. To enable a comparison with them, we captured several demos from their project websites and conducted tests using the same target face image. The comparative results are shown in the Fig. 14. Our approach better preserves the original video attributes such as lighting and expression, and also demonstrates strong stability in occluded cases.
Appendix K Limitations
LivingSwap achieves better fidelity by directly referencing the source video, while enabling more flexible identity control and improved stability in long videos through keyframe identity injection. However, this framework design also introduces certain limitations: 1) Dependence on Keyframe Quality: The keyframe identity injection method creates a reliance on the quality of the selected keyframes. Although we demonstrated in Appendix C that LivingSwap shows strong robustness to keyframe quality, it still encounters issues when dealing with keyframes that yield poor face-swapping results (e.g., identity or expression drift, image distortion, etc.). In such cases, the final output can be biased by the keyframe. This is supported by the experimental results in Appendix D, where selecting higher-quality keyframes led to significant improvements on the 10 worst-performing cases. 2) Slow Inference Speed: The use of a large video dataset to train the DiT model, combined with the Attribute Encoder for condition injection, significantly impacts the face-swapping speed of LivingSwap. In our experiments, for a video of 81 frames (approximately 3 seconds at 25 fps), LivingSwap requires 195 seconds (about 3 minutes) on a single H200 GPU with 108 GB memory. This translates to approximately 1 minute per second of swapped video. In conclusion, our future work will focus on exploring better methods for identity injection and optimizing the face-swapping speed.